OpenStudy (frank0520):

Can anyone help me with this complex vector question:

OpenStudy (frank0520):

here is the question:

OpenStudy (irishboy123):

@Empty @ganeshie8

OpenStudy (irishboy123):

@jango_IN_DTOWN

OpenStudy (amoodarya):

can you find eigenvalues ?

OpenStudy (frank0520):

So I found: \[P=\left[\begin{matrix}-1/2 & -1/2 \\ 1 & 0\end{matrix}\right]\] \[P^{-1}=\left[\begin{matrix}0 & 1 \\ -2 & -1\end{matrix}\right]\] and \[\alpha=\sqrt{17}\] I need help finding the angle.

OpenStudy (amoodarya):

1+4i 1-4i ?

OpenStudy (frank0520):

for eigen values I had: -1+4i -1-4i

OpenStudy (frank0520):

no, can you explain that?

OpenStudy (amoodarya):

\[\left[\begin{matrix}-1-4i & 0 \\ 0 & -1+4i\end{matrix}\right]\] this is diagonal jordan form of you matrix

OpenStudy (frank0520):

yes

OpenStudy (amoodarya):

now multiply S and R and then solve SR=J then you will find cos theta ,...

OpenStudy (amoodarya):

sorry my net is down . ...